### v1

Abstract
The dimensional transition in turbulent jets of a shear-thinning fluid is studied via direct numerical simulations. Our findings reveal that under vertical confinement, the flow exhibits a unique mixed-dimensional (or $2.5$D) state, where large-scale two-dimensional and small-scale three-dimensional structures coexist. This transition from three-dimensional turbulence near the inlet to two-dimensional dynamics downstream is dictated by the level of confinement: weak confinement guarantees turbulence to remain three-dimensional, whereas strong confinement forces the transition to two-dimensions; the mixed-dimensional state is observed for moderate confinement and it emerges as soon as flow scales are larger than the vertical length. In this scenario, we observed that the mixed-dimensional state is an overall more energetic state and it shows a multi-cascade process, where the direct cascade of energy at small scales and the direct cascade of enstrophy at large scales coexist. The results provide insights into the complex dynamics of confined turbulent flows, relevant in both natural and industrial settings.
keywords:
Jets, shear-layer turbulence, turbulence simulation
1 Introduction
The flow of a low-viscosity fluid at high-speed is chaotic in nature. The energy injected to sustain this state is transferred from large to small eddies, down to a particular scale from which it is dissipated by the viscosity of the fluid (Kolmogorov, 1941). Conventional turbulence in three-dimensions fulfills this description, whereas new phenomena appear in two-dimensions. Energy transfer in two-dimensional turbulence is dictated by a double cascade scenario: an inverse cascade of kinetic energy to large scales and a direct cascade of enstrophy (squared vorticity) to small scales (Kraichnan, 1967; Batchelor, 1969; Boffetta, 2007). Certainly, no physical system is two-dimensional in reality, though two-dimensional turbulence becomes relevant if one spatial direction is greatly constrained, e.g., by geometry (Boffetta et al., 2012; Boffetta & Ecke, 2012). For instance, the large-scale motions in atmospheric flows comply with two-dimensional turbulence (Charney, 1971; Nastrom et al., 1984; Lindborg, 1999). In this case, the flow domain is subdued to a large aspect ratio: the horizontal lengths are much larger than the height of the atmospheric layer.
The confinement in thin layers can induce a rich phenomenology in turbulent flows that, if forced at intermediate scales, produces a split energy cascade (Smith et al., 1996; Celani et al., 2010; Alexakis & Biferale, 2018). Under this circumstance, a portion of the energy flows to the large scales in a two-dimensional-fashion. Conversely, the remaining part cascades toward the small viscous scales. Interestingly, a direct cascade of enstrophy can develop simultaneously at scales smaller than the forcing but larger than the thickness of the layer, and three-dimensionality becomes relevant only at much smaller scales (Musacchio & Boffetta, 2017). Nevertheless, the presence of physical confinement, e.g., using solid boundaries, is not compulsory to observe the split energy cascade. In fact, numerical simulations in a fully-periodic box with one dimension much smaller than the others have shown this phenomena (Smith et al., 1996; Celani et al., 2010). Despite this, its occurrence changes with the boundary conditions. For example, the development of the shear layer in wall-bounded flows restricts the development of two-dimensional dynamics (Xia et al., 2011; Byrne et al., 2011; Boffetta et al., 2023).
Here, we consider a planar jet, i.e., the flow is injected through a plane slit of half-width $h$ in a computational box periodic in the vertical direction $z$. We adopt a shear-thinning fluid in which the viscosity decreases non-linearly for increasing values of the shear rate. This setup is representative of the outflow of a river into the sea or the ocean, in which a stream of freshwater rich in bacteria or microalgae is delivered into sea water. Indeed, the presence of bacteria and microalgae in suspension has been shown to introduce non-Newtonian features into the fluid, such as shear-dependent viscosity (Al-Asheh et al., 2002; Zhang et al., 2013). Differences in salinity, temperature and density between the freshwater stream and the salt water can impede mixing, thus leading to the formation of a stratified flow with a (thin) layer of freshwater flowing over salt water. The freshwater layer rich in microalgae and bacteria is reminiscent of our computational setup, where the height of the domain (the thickness of the freshwater layer) is much smaller than the other two dimensions.
Within this framework, we show that when the thickness of the domain is large, the flow is completely three-dimensional, while when it is small, it is fully two-dimensional. Interestingly, for intermediate cases, the flow spatially transitions from $3$D close to the inlet to $2$D further downstream, with the two regimes being connected by a region of mixed-dimensional turbulent dynamics where the constraint modulates the largest scales towards two-dimensions and the smaller ones remain three-dimensional. We indeed observe, at intermediate levels of vertical constraint, a multi-cascade process, where both a direct cascade of energy at small scales and a direct cascade of enstrophy at large scales coexist.
2 Setup and method
2.1 Problem setup
We have addressed this study by means of three-dimensional direct numerical simulations. The planar jet is generated by fluid injected with a uniform velocity $U$ through a planar slit spanning the entire height of the domain, and it develops over a stream-wise distance $L_{x}=160h$ in a computational box of width $L_{y}=240h$ and height $L_{z}$. The vertical length $L_{z}$ is varied among simulations: we consider five distinct simulations with $L_{z}=0.83h,1.67h,3.33h,6.67h,13.33h$, respectively. The thinnest domain ($L_{z}=0.83h$) introduces a strong vertical constraint in order to allow the development of a two-dimensional flow, which is progressively relaxed as $L_{z}$ is increased while maintaining a thin computational domain, so $L_{z}\ll L_{x}\sim L_{y}$.
The flow dynamics are simulated by solving the Navier-Stokes equations:
$\nabla\cdot\boldsymbol{u}=0,$ (1)
$\rho\left(\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla{\mathbf{u}}\right)=-\nabla p+\nabla\left[\mu\left(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}\right)\right],$ (2)
where $\mathbf{u}$ is the local flow velocity, $\rho$ the density, $p$ the pressure and $\mu$ the local viscosity. We adopt an inelastic, shear-thinning fluid, whose behaviour is defined via the Carreau fluid model (Bird et al., 1974). The local viscosity $\mu$ depends on the local shear rate $\dot{\gamma}$ as:
$\mu(\dot{\gamma})=\mu_{\infty}+\left(\mu_{0}+\mu_{\infty}\right)\left[1+\left(\lambda\dot{\gamma}\right)^{2}\right]^{\frac{n-1}{2}},$ (3)
where $\lambda$ is the fluid consistency index, and $\mu_{0}$ and $\mu_{\infty}$ are the zero-shear viscosity and the viscosity for $\dot{\gamma}\rightarrow\infty$, respectively. We set the power-law index $n=0.2$, thereby obtaining a strong shear-thinning effect. We select a low value of the inlet Reynolds number (ratio of inertial to viscous effects), $Re=\rho hU/\mu_{0}=20$. It should be noted that Newtonian planar jets are laminar at this value of $Re$ (Deo et al., 2008; Sureshkumar & Beris, 1995; Sato & Sakao, 1964; Soligo & Rosti, 2023), thus any turbulent motion is caused exclusively by the shear-thinning in the flow. Turbulence however is still Newtonian, as it originates by the prevalence of inertial over viscous terms. The non-Newtonian character of the flow promotes indeed the onset of the instability, so the transition to turbulence is at markedly lower $Re$ compared to Newtonian planar jets (Ray & Zaki, 2015; Yamani et al., 2023; Soligo & Rosti, 2023). The non-Newtonian contribution is described using the Carreau number, defined as $Cu=h\lambda/U=100$.
Our numerical setup is easily testable by experiments (recent experimental work by Yamani et al. (2023) addressed the flow of a viscoelastic planar jet at low Reynolds), whereas the effect of physical confinement is attenuated by the shear-thinning characteristic of the fluid, which reduces the extent of the shear layer (viscosity decreases at the wall boundaries in the experimental setup).
2.2 Numerical method
The motion of the incompressible, shear-thinning fluid is solved by simulating the Navier-Stokes equations with variable viscosity using the in-house solver Fujin (https://groups.oist.jp/cffu/code). The Navier-Stokes equations are discretized on a uniform, staggered, Cartesian grid; the fluid velocities are located at the cell faces, whereas pressure and viscosity at the cell centers. The fluid viscosity is updated at every time-step following Eq. 3. The spatial derivatives are approximated using second-order finite differences in all directions. The system is advanced in time through a second-order Adams-Bashforth scheme coupled with a fractional step method (Kim & Moin, 1985). The divergence-free velocity field is enforced by solving the Poisson equation for the pressure using an efficient solver based on the fast Fourier transform. We resort to the domain decomposition library 2decomp (http://www.2decomp.org) and the MPI protocol to parallelize the solver.
We adopt a uniform grid spacing in all spatial directions, in $x$ and $y$ the domain is discretized using $N_{x}\times N_{y}=1536\times 2304$ grid points, while the number of points in the $z$ direction depends on the height of the domain, namely $N_{z}=8,16,32,64,128$ for increasing heights. The jet fluid is issued through a slit at $x=0$; at this boundary we impose no-slip and no-penetration boundary conditions, exception made for the inlet part, where we impose a plug-flow velocity profile. At the outlet boundary ($x=L_{x}$) we use a non-reflective boundary condition (Orlanski, 1976). At the side boundaries ($y=0$ and $y=L_{y}$) we impose free-slip and no penetration boundary conditions. Lastly, at the top and bottom boundaries ($z=0$ and $z=L_{z}$) we impose periodic boundary conditions.
3 Results
Figure 1: Effect of the constrained dimension $L_{z}$ in the turbulent planar jets. Top panel: magnitude of the instantaneous vorticity field $\left|\left|\boldsymbol{\omega}\right|\right|U^{-1}h$ in the $z=L_{z}/2$ plane for the three-dimensional, mixed-dimensional and two-dimensional planar jets, from left to right. The inset shows a zoomed view of the region near the inlet at $x=\left[0,20h\right]$, $y=\left[-5h,5h\right]$. Bottom panel: power spectra of the turbulent kinetic energy computed at the jet centerline at distances $x=40h,80h,120h$ from the inlet. The spectra follows the typical $f^{-5/3}$ scaling if turbulence is primarily three-dimensional (red). In two-dimensional turbulence (blue), instead, the flow exhibits a $f^{-3}$ decay that becomes steeper as $x$ increases, tending towards $f^{-5}$ in the most constrained planar jet.
Figure 1 shows the impact of $L_{z}$ on the turbulent planar jets, in which, as anticipated, the constraint dictates the turbulent regime in the flow. Consequently, the morphology of the vorticity structures changes significantly with $L_{z}$ (see top panel in Fig. 1). We first observe that turbulence is three-dimensional if the flow is not constrained, i.e., for sufficiently large values of $L_{z}$. A complete different flow structure is instead observed at low values of $L_{z}$, in which large vortices form dipoles (pairs of counter-rotating vortices) that are advected downstream, and no small scale flow structures are observed, thus indicating that turbulence is mainly two-dimensional. The flow does not transition in bulk from $3$D to $2$D when changing $L_{z}$, with the planar jet experiencing an intermediate state where both three-dimensional and two-dimensional structures are present in the flow at the same time. This transitional regime, hereafter termed mixed-dimensional (or $2.5$D), is characterized by the simultaneous coexistence in the flow of large-scale two-dimensional and small-scale three-dimensional structures.
Next, we inspect the energy spectra in the bottom panel of Fig. 1. Note that we compute the velocity spectra in time rather than in space by recording velocity data over time from a probe placed on the centerline of the jet, similarly to what done in experiments. Computing the power spectra in time rather than in the vertical direction allows us to have a wider energy spectrum that is not limited by the height of the domain. The equivalence of time and space spectra has been demonstrated in the past (Namer & Ötügen, 1988; Soligo & Rosti, 2023). As clearly shown in the figure, $L_{z}$ influences the energy cascade, which depicts a different behavior depending on the regime of turbulence. First, the least constrained jet shows the conventional $f^{-5/3}$ scaling for three-dimensional turbulence (Kolmogorov, 1941). Consistently, we observe the energy cascade typical of three-dimensional turbulence: the jet instability gives energy to the flow and it generates large structures that break down in progressively smaller and smaller eddies. As the vortices move downstream, the characteristic shear rate reduces – hence viscosity increases – and energy is dissipated. We expect to recover the power-law scaling for the three-dimensional turbulence energy cascade as the turbulent motions are Newtonian: they are generated by the competition of inertial and viscous terms (Soligo & Rosti, 2023). Eventually, dissipation becomes relevant at every scale and the cascade is impeded: the scaling $f^{-5/3}$ is not present at $x=120h$. On the other hand, the most constrained case exhibit two-dimensional flow and a scaling of $f^{-3}$. Here, the three-dimensional cascade is clearly disrupted and two-dimensional phenomena become dominant (Kraichnan, 1967; Batchelor, 1969). The change in the flowing regime observed here is due only to the vertical confinement: we adopt a non-Newtonian fluid model which is characterized by the presence of shear-thinning alone (there are no viscoelastic effects). The spectrum becomes steeper as $x$ increases and it eventually seems to saturate at $f^{-5}$. The steepening of the energy spectrum agrees with the appearance of dispersed large-size coherent vortices in the flow (Basdevant et al., 1981; McWilliams, 1984; Benzi et al., 1986; Legras et al., 1988). Note that the change in the vertical constraint also alters the instability in the region close to the inlet: for strong vertical constraints (small $L_{z}$) we observe a flapping motion of the shear layers, whereas puffing motion dominates when the constraint is relaxed (large $L_{z}$). The flapping dynamic is associated with the antisymmetric, sinuous mode, that destabilizes the flow more substantially, thus injecting more energy (Mattingly & Criminale Jr, 1971), as also observed by the energy spectra close to the inlet which are shifted upwards in 2D compared to the 3D case. At last, the mixed-dimensional planar jet exhibits features from both three- and two- dimensional planar jets: the scaling $f^{-5/3}$ is found close to the inlet and it changes towards $f^{-3}$ downstream. The spectrum becomes slightly steeper further downstream, consistent with the appearance of the coherent vortices. Similarly to the two-dimensional jet, the flow is more energetic close to the inlet. Very close to the inlet, the vorticity fluctuations look closely related to those in the three-dimensional case, implying the existence of puffing events connected to the varicose mode (Sato, 1960; Mattingly & Criminale Jr, 1971). However, flapping motions become soon dominant and the macroscopic vorticity structures resemble those from the two-dimensional planar jet, thus indicating the presence of the more energetic sinuous mode.
Figure 2: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations for different vertical constraints $L_{z}$. We compute $S_{3}$ at $x=40h$ at the flow centerline for each planar jet: $L_{z}=0.83h,1.67h,3.33h,6.67h,13.33h$. We show the absolute value $\left|S_{3}\right|$ for clarity, and we indicate with colors whether $S_{3}$ is positive (blue) or negative (red), so that the corresponding turbulent scales are either two- or three- dimensional. Consequently, the scaling for $\left|S_{3}\right|$ changes as $L_{z}$ is shifted towards larger values. We identify two scalings: $S_{3}\sim r^{3}$ if turbulence is two-dimensional, corresponding to a direct cascade of enstrophy, and $S_{3}\sim-r$ if it is three-dimensional, thus denoting a direct cascade of energy. The plots are shifted vertically for better readability.
To better characterise the different nature of the turbulent fluctuations at each scale, we now calculate the longitudinal velocity differences $\Delta u(r)=\left(\boldsymbol{u}\left(\boldsymbol{x}+\boldsymbol{r}\right)-\boldsymbol{u}\left(\boldsymbol{x}\right)\right)\cdot\boldsymbol{r}/\left|\boldsymbol{r}\right|$. Concretely, we introduce the third-order structure function $S_{3}(r)=\langle\left(\Delta u\right)^{3}\rangle$, where the angle brackets indicate averaging in time and in space, shown in Fig. 2 for the different cases analysed. A remarkable property of $S_{3}$ is that it denotes whether the flow scales are two- or three- dimensional depending on its sign (Lindborg, 1999; Kolmogorov, 1991), and it can help understanding the direction of the energy and enstrophy fluxes (Cho & Lindborg, 2001; Bernard, 1999; Cerbus & Chakraborty, 2017). In the particular case of two-dimensional turbulence, $S_{3}$ is positive (Lindborg, 1999) whereas it is negative if turbulence is three-dimensional (Kolmogorov, 1991). Indeed, we observe that $S_{3}$ is positive (two-dimensional flow) for the most constrained case ($L_{z}=0.83h$) whereas is is negative (three-dimensional flow) for the least constrained case ($L_{z}=6.67h,13.33h$). Furthermore, the scaling of $S_{3}$ outlines the preferred cascade process, thus indicating the direct cascade of enstrophy in the two-dimensional case ($S_{3}\sim r^{3}$) and the direct cascade of energy in the three-dimensional counterpart ($S_{3}\sim-r$). We do not find evidence of an inverse mechanism of energy transfer toward the large scales in the two-dimensional case. This is not surprising if we consider that energy is injected in the flow through the planar slit, thus not forcing the flow at any intermediate scale (Boffetta, 2007; Boffetta & Ecke, 2012). More interestingly, turbulence is characterized by a mixed-dimensional regime for the right choice of $L_{z}$. Two- and three- dimensional scales are present simultaneously in the flow for $L_{z}=1.67h,3.33h$. In these cases, large scales are two-dimensional while the small ones are three-dimensional, with the transition between regimes happening at some intermediate scale. This transition is strongly dependent on the height of the domain and it can be delayed further in $x$ for increasing values of $L_{z}$.
Figure 3: Third-order structure functions $S_{3}$ of the longitudinal velocity fluctuations for different distances from the inlet. Panel a) shows $S_{3}$ computed at distances $x$ uniformly separated along the jet centerline ($x=20h,30h,\ldots,140h$) for the mixed-dimensional planar jet with $L_{z}=3.33h$. We indicate whether the turbulent scales are two- or three- dimensional with colors, similarly to Fig. 2. The plots are shifted vertically for better readability. The inset shows the scale $r$ at which $S_{3}$ changes its sign. Red and blue markers are used for fully $3$D and $2$D structure functions respectively, while purple markers for those showing both regimes. Note that, if $S_{3}$ does not change sign, we mark either the largest three-dimensional scale or the smallest two-dimensional one. Panel b) depicts $S_{3}$ for the least constrained ($L_{z}=13.33h$) or $3$D jet, and panel c) for the most constrained ($L_{z}=0.83h$) or $2$D jet. $S_{3}$ is plotted in a similar fashion than a), where we computed $S_{3}$ at three distances $x=40h,80h,120h$ at the jet centerline. In these cases, $S_{3}$ has the same sign throughout $x$, thus observing a single scaling at all distances dependent on the dimensionality of the flow.
To further investigate the transition between regimes, we choose a mixed-dimensional planar jet, in particular $L_{z}=3.33h$, and we calculate repeatedly $S_{3}$ at several distances from the inlet. Results are summarized in Fig. 3a. Close to the inlet the flow is three-dimensional: the effect of the vertical constraint does not hinder the development of the three-dimensional regime. The largest flow scale is smaller than the geometrical constraint, thus not compromising the energy cascade of the three-dimensional flow. We report the presence of a single scaling, $S_{3}\sim-r$, typical of the direct energy cascade. At intermediate distances, $x=\left[40h,100h\right]$, a mixed-dimensional regime appears: both two- and three-dimensional regimes coexist, with three-dimensional turbulence characterizing the small scales and two-dimensional turbulence characterizing instead the large scales. Here, the vertical confinement hinders the development of three-dimensional turbulence at the largest scales, while it has no effect on the three dimensional turbulence at the smallest scales. In this case, we have the simultaneous presence of a direct energy cascade ($S_{3}\sim-r$) at small scales and a direct enstrophy cascade ($S_{3}\sim r^{3}$) at large scales. We report in the inset of Fig. 3a the scale at which the structure function changes sign, i.e., the scale where the flow transitions from $3$D to $2$D. We observe that, as soon as the jet thickness reaches the size of the vertical confinement, the largest scales become two-dimensional. However, some scales between the largest ones and the vertical confinement remain three-dimensional, causing the largest $3$D structures to be anisotropic. Moving downstream, the anisotropy of the largest $3$D structures grows till a value of around $3L_{z}$ at $x\approx 60h$, after which the flow becomes more and more two-dimensional. Eventually, we observe that the flow becomes completely two-dimensional at the farthest distances from the inlet, $x\geq 120h$, where the characteristic flow scales are the largest and the vertical confinement impedes the development of three-dimensional flow at all scales. We indeed observe a scaling of the structure function $S_{3}$ compatible with the direct enstrophy cascade ($S_{3}\sim r^{3}$). The range of scales observed at each distance from the inlet depends on two factors: the characteristic length scale of the jet (the jet thickness), and the local viscosity. Both of these quantities increase with increasing distance from the inlet of the jet; the jet thickness determines the largest scale in the flow, while the local viscosity is among the factors determining the eventual development of three-dimensional turbulent motions.
Lastly, we perform the same analysis on the the least ($L_{z}=13.33h$) and the most ($L_{z}=0.83$) constrained jets at several locations on the jet centerline ($x=40h,80h,120h$) (see Fig. 3b, c). In these cases, $S_{3}$ has the same sign throughout $x$, indicating that turbulence is either three- ($S_{3}<0$) or two- ($S_{3}>0$) dimensional in the planar jet. The structure functions show power-law scaling for intermediate values of the separation distance $r$ in the three-dimensional planar jet; the scaling law $S_{3}\sim-r$ agrees with the direct cascade of energy. On the other hand, the scaling for $S_{3}$ in the two-dimensional planar jet shifts towards $S_{3}\sim r^{3}$, indicating instead the presence of the direct cascade of enstrophy. More interestingly, the scaling holds for small separation distances farther from the inlet. This is in good agreement with the concept of the enstrophy cascade as a space-filling phenomenon, thus being present at very small scales (Benzi et al., 1986). Additionally, $S_{3}$ follows an anomalous behavior at large and intermediate $r$. As observed in Fig. 1, large-size coherent vortices emerge far away from the inlet, which interrupt the enstrophy cascade. At these distances, turbulence becomes more intermittent as velocity fluctuations are localized in the vortexes.
4 Conclusions
We have studied via direct numerical simulations how the vertical confinement of a turbulent planar jet can alter its dimensionality. We show that, under the right constraint, a mixed-dimensional turbulent regime appears, that is characterized by the simultaneous presence of large-size two-dimensional and small-size three-dimensional scales. The onset of this particular regime is dictated by the size of the constrained dimension $L_{z}$: as soon as the flow scales become larger than $L_{z}$, the flow dimensionality changes from three- to two- dimensions. This transition is postponed further downstream as the flow is less constrained (increasing $L_{z}$). Therefore, for sufficiently large $L_{z}$, turbulent scales are simply three-dimensional along the jet, and the direct cascade of energy is enabled. Conversely, a strong confinement (small $L_{z}$) makes the flow two-dimensional: the direct cascade of energy is disrupted and the direct cascade of enstrophy takes place. Both cascades are conserved wherever the mixed-dimensional turbulent state is present: the direct cascade of energy is active at small scales, whereas the direct cascade of enstrophy dominates at large scales. In addition, we found that the mixed-dimensional configuration is an overall more energetic state, thus partially retaining the three-dimensionality in the flow while deferring the emergence of two-dimensional strong vortical structures further downstream from the inlet. The direct enstrophy and energy cascades and the respective scalings we report here are the same of a Newtonian fluid: the Carreau fluid is a non-Newtonian fluid model characterized by shear-thinning alone, which allows to attain Newtonian turbulence at a relatively low Reynolds number.
[…]
[Acknowledgements]The research was supported by the Okinawa Institute of Science and Technology Graduate University (OIST) with subsidy funding from the Cabinet Office, Government of Japan. The authors acknowledge the computational resources provided by the Scientific Computing section of the Research Support Division at OIST.
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
[Author ORCIDs]
Christian Amor, https://orcid.org/$0$-0002-9710-7917
Giovanni Soligo, https://orcid.org/$0$-0002-0203-6934
Andrea Mazzino, https://orcid.org/$0$-0003-0170-2891
Marco Edoardo Rosti, https://orcid.org/$0$-0002-9004-2292